Suppose that the real money demand function is 𝐿(𝑌, 𝑟 + 𝜋 𝑒 = 0.3𝑌 𝑟 + 𝜋 𝑒 Where Y is real output, r is the real interest rate, and πe is the expected rate of inflation. Real output is constant over time at Y = 1500. The real interest rate is fixed in the goods market at r = 0.5 per year. (a) Suppose that the nominal money supply is growing at the rate of 10% per year and that this growth rate is expected to persist for ever. Currently, the nominal money supply is M = 400. What are the values of the real money supply and the current price level? (Hint: What is the value of the expected inflation rate that enters the money demand function?). (10) (b) Suppose that the nominal money supply is M = 400. The Bank of Namibia announces that from now on the nominal money supply will grow at the rate of 5% per year. If everyone believes this announcement, and if all markets are in equilibrium, what are the values of real money supply and the current price level? (10) .


1
Expert's answer
2021-04-28T07:33:41-0400

(a) inflation rate=growth rate of money supply-growth rate of output

"0.1-0.005=0.095\u00d7100=9.5" %

Growth rate of supply+growth rate of money=inflation rate+growth rate of output

"0.1+v=0.095+0.005"

"0.1+v=0.1"

"V=0"

Current price level=(nominal money supply×velocity of money )/real output

"400\u00f70=0"

Real money supply=nominal money supply/price level

"400\u00f70=400"


(b)inflation rate=growth rate of money supply-growth rate of output

"0.05-0.005=0.045\u00d7100=4.5" %

Growth rate of supply+growth rate of money=inflation rate+growth rate of output

"0.05-v=0.045+0.005"

"V=0"

Current price level=(nominal money supply×velocity of money )/real output

"4000\u00f70=0"

Real money supply=nominal money supply/price level

"400\u00f70=400"